# VecCity: A Taxonomy-guided Library for Map Entity Representation Learning [Experiment, Analysis & Benchmark]

### v1

@@ 4.2.2. Experimental Results for Road Segment MapRL
2) Multiple Pre-training Tasks Improve Performance: Models with multiple types of pre-training tasks, i.e., Toast, HRoad, START, and JCLRNT, outperform those with single-task pre-training, i.e., HRNR and SARN. Notably, among token-based + graph-based models, HRoad, which integrates both NFI and GAu tasks, surpasses HRNR (GAu only) and SARN (AGCL only). Using diverse pre-training tasks enables the model to extract information from multiple perspectives, enhancing its ability to learn more comprehensive feature representations. Moreover, in the full-pipeline models, JCLRNT combines a sequence-oriented task (Augmented Trajectory Contrastive Learning) with a graph-oriented task (Augmented Graph Contrastive Learning) and outperforms START, which uses two sequence-oriented tasks (MTR and ATrCL). This suggests that even in models with multiple tasks, greater task heterogeneity can further enhance performance.
3) Efficiency Trade-offs: Simpler models with fewer modules, such as SARN and RN2Vec, are significantly more efficient. Models without sequence encoders, such as HRNR and HRoad, also exhibit better efficiency since incorporating trajectory data demands considerable memory and time. Improving the efficiency of trajectory-based models remains a key challenge for future research.
Dataset Beijing Chengdu Xi’an Porto San Francisco Overall
Task LPC FI LPC FI LPC FI LPC FI LPC FI Per Ave Rank Per Rank
Metric ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓ ACC@1↑ F1↑ MAE↓ RMSE↓
ZEMob 0.380 0.110 30.90 55.19 0.339 0.115 154.9 441.2 0.316 0.138 133.7 326.2 0.298 0.204 527.5 595.1 0.820 0.477 509.1 910.5 5.2 7
GMEL 0.412 0.571 27.68 58.91 0.338 0.109 154.8 441.1 0.330 0.359 133.1 324.4 0.267 0.335 525.1 591.9 0.620 0.550 514.0 916.7 4.5 6
MGFN 0.567 0.601 30.09 54.45 0.345 0.135 154.6 448.2 0.330 0.162 132.7 324.9 0.304 0.246 531.3 596.9 0.835 0.533 605.3 1011 4.4 5
HDGE 0.409 0.229 29.26 54.82 0.365 0.157 159.6 402.1 0.473 0.242 134.1 334.5 0.327 0.263 525.2 576.9 0.820 0.450 511.8 926.4 4.3 4
MVURE 0.479 0.267 28.13 51.37 0.485 0.279 161.5 450.2 0.545 0.235 133.7 326.0 0.316 0.294 524.8 592.4 0.820 0.450 511.2 913.1 4.1 3
ReMVC 0.722 0.686 27.81 52.66 0.610 0.542 147.7 430.9 0.567 0.412 131.5 324.6 0.404 0.357 527.3 594.8 0.820 0.450 444.7 852.3 3.0 2
HREP 0.524 0.405 22.91 45.85 0.528 0.445 124.0 309.6 0.611 0.494 127.9 312.3 0.386 0.358 498.0 574.5 0.836 0.538 438.4 799.3 2.2 1
Task MI Efficiency MI Efficiency MI Efficiency MI Efficiency MI Efficiency Eff Ave Rank Eff Rank
Metric MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓ MAE↓ RMSE↓ Param↓ Speed↓
ZEMob 0.379 0.771 45.10 40 3.599 29.66 7.699 19 3.776 17.63 5.670 1 44.85 72.30 1.047 1 5.837 23.48 2.335 1 4.0 3
GMEL 0.418 0.712 0.118 713 4.599 29.66 0.118 193 3.668 17.60 0.118 189 41.08 57.41 0.118 161 7.377 22.29 0.118 2 4.0 3
MGFN 0.283 0.694 1156 839 5.176 29.23 485.7 6 3.889 17.23 318 2 39.37 56.16 53.67 2 6.846 22.22 42.19 1 5.4 7
HDGE 0.389 0.759 12.91 1799 5.189 29.00 1.505 601 3.863 16.85 1.217 75 38.69 54.20 0.440 117 7.397 21.94 0.223 3 5.1 6
MVURE 0.281 0.685 0.097 12 5.200 29.37 0.097 8 3.941 17.41 0.097 1 38.63 62.21 0.097 1 7.196 22.73 0.097 1 2.1 1
ReMVC 0.276 0.667 1.450 873 5.230 29.42 0.181 582 3.986 17.40 0.149 39 40.34 56.66 0.069 47 7.007 22.60 0.050 7 4.1 5
HREP 0.318 0.677 0.182 9 5.270 28.33 0.182 7 3.949 17.05 0.182 1 37.42 57.62 0.182 2 6.754 21.71 0.182 1 3.3 2
@@ 4.2.3. Experimental Results for Land Parcel MapRL
Table 5 presents the performance comparison of land parcel MapRL models on the downstream tasks of Land Parcel Classification (LPC), Flow Inference (FI), and Mobility Inference (MI). The following observations can be made:
@@
3) Need for More Diverse Pre-training Tasks: Current models mainly rely on the Graph Autoencoder (GAu) task for pre-training. However, experimental results show that tasks such as Augmented Token Contrastive Learning (AToCL) and Neighborhood Contrastive Learning (NCL) yield better performance. Developing models that incorporate a broader range of pre-training tasks is a promising direction for enhancing land parcel MapRL models.
@@ 4.2.4. Summary for Overall Experiments
3) Role of Auxiliary Data: Incorporating POI, trajectory, and relation network data improves model performance across tasks. However, models that fully exploit trajectory dependencies achieve diminishing gains from relation networks, indicating some redundancy between these data sources.
4) Trade-offs Between Performance and Efficiency: While complex models with multiple encoders and tasks offer higher accuracy, they require more computational resources. Simpler token-based models remain attractive for applications where efficiency is critical, despite offering slightly lower performance.
4.3. Performance over Small Datasets
In the overall experiments, both the pre-training and fine-tuning steps are executed on the full training dataset. However, in practical applications, MapRL models are often trained on abundant data while fine-tuned using limited labeled samples for downstream tasks. To assess the robustness of mainstream models under such conditions, we conduct a small dataset experiment.
In this experiment, we reduce the fine-tuning data size from 50% to 10% of the original training set and compare the result of pre-built models in VecCity. POI-oriented models are evaluated on the Tokyo dataset, while road-segment and land-parcel-oriented models are tested on the Xi’an dataset. Figure 5 presents the evaluation result: Fig. 5 (a) to (c) show results for POI-oriented models. Fig. 5 (d) to (f) display results for road-segment-oriented models. Fig. 5 (g) to (i) provide results for land-parcel-oriented models. From these figures, we make following observations:
@@ 4.3. Performance over Small Datasets
1) Stability in Attribute Inference Tasks: MapRL models show stable performance across varying training data sizes in attribute inference tasks, such as POI Classification (POIC) in Fig. 5 (a), Average Speed Inference (ASI) in Fig. 5 (d), and Land Parcel Classification (LPC) in Fig. 5 (g). This stability arises because these tasks focus on inferring entity attributes, which are effectively captured during pre-training. As a result, the fine-tuning phase requires only a small amount of labeled data to maintain good performance.
2) Fluctuations in Trajectory-Related Tasks: Models exhibit significant performance fluctuations in trajectory-related tasks, such as Next POI Prediction (NPP) in Fig. 5 (b), Trajectory User Link (TUL) in Fig. 5 (c), and Travel Time Estimation (TTE) in Fig. 5 (e). These tasks rely on sequence-level predictions, while encoders and pre-training tasks focus on generating single-vector representations for individual entities. This mismatch makes trajectory-related tasks more sensitive to the size of the fine-tuning dataset, as limited labels reduce the model’s ability to learn sequential dependencies effectively. For Similarity Trajectory Search (STS) in Fig. 5 (f), although it is a trajectory-related task, it involves similarity comparison rather than prediction so is less affected by this mismatch.
3) Fluctuations in Flow and Mobility Inference Tasks: Models also show significant performance fluctuations in Flow Inference (FI) in Fig. 5 (h) and Mobility Inference (MI) in Fig. 5 (i), even though these tasks are not directly related to trajectory data. This is because the labels for these tasks reflect statistical data on human mobility flows over specific periods, which are highly time-heterogeneous, with label distributions varying significantly across periods. As a result, more labeled data is required to capture these complex temporal patterns accurately.
@@
Overall, existing MapRL models exhibit varying sensitivity to the size of labeled data across different tasks. For tasks less influenced by temporal information, such as attribute inference (ASI, POIC, LPC) and Similarity Trajectory Search (STS), models maintain stable performance even with limited training data. However, for tasks with time-dependent labels, such as trajectory-related tasks (NPP, TUL, TTE) and flow/mobility inference tasks (FI, MI), larger amounts of labeled data are necessary to effectively capture temporal dynamics.
Table 6. Results for CTLE variants on the Tokyo dataset.
Methods CTLE CTLECTLE& CTLE* CTLE\$
Encoder Core Arc TE TE TE TE LSTM
Pretraining Task TrajP
MTR
ATrCL
POIC ACC@1↑ 0.048 0.032 0.041 0.050 0.043
F1↑ 0.006 0.002 0.005 0.007 0.006
NPP ACC@1 0.154 0.089 0.153 0.165 0.058
ACC@5↑ 0.336 0.193 0.332 0.352 0.147
TUL ACC@1 0.698 0.401 0.687 0.699 0.388
F1↑ 0.444 0.160 0.437 0.448 0.104
Per Rank 2 5 3 1 4
4.4. Comparing of Modified Models
In this section, we conduct experiments to evaluate the influence of pre-training tasks and encoder structures on the performance of downstream tasks. We modify three representative MapRL models – CLTE, JCLRNT, and HREP – which achieve the highest average performance in POI, road segment, and land parcel representation learning, respectively. Owing to VecCity’s modular design, specific pre-training tasks can be easily removed from models, and new pre-training tasks can be seamlessly integrated.
The experimental results are shown in Tab. 68. In the figures, the black “✓” and “✗” indicate whether the corresponding pre-training tasks were originally included or excluded when the model was first proposed. In contrast, the red “” and “” indicate newly added or removed pre-training tasks. Additionally, we evaluate the impact of different backbone models by modifying the core architecture of encoders (highlighted in red in the “Encoder Core Arc” row). In the last row, we present the performance ranking of each variant for easier comparison and conclusion.
Comparison of CTLE Variants. Among POI MapRL models, CTLE achieves the best performance. In Tab. 6, we experiment with the following model variants for CTLE. (a) CTLE: Maintains the original model architecture. (b) CTLE%: Removes the Masked Trajectory Recovery (MTR) pre-training task. (c) CTLE&: Removes the Trajectory Prediction (TrajP) pre-training task. (d) CTLE*: Integrates a new Augmented Trajectory Contrastive Learning (ATrCL) pre-training task, using the ATrCL task proposed by CACSR, which achieved the second-best performance in the overall POI experiments. (e) CTLE\$: Replaces the core architecture of CTLE (a Transformer encoder) with an LSTM layer.
The downstream task results are presented in Tab. 6. From these experiments, we draw the following observations: $i$) New Task Enhances Performance: The CTLE* variant achieves the best performance across all downstream tasks, indicating that integrating the new ATrCL task significantly improves the model’s capability. $ii$) MTR Provides Richer Representations: The CTLE& variant outperforms CTLE%, suggesting that the MTR task offers richer representation capabilities than the TrajP task. $iii$) Transformer Encoder is Essential for Trajectory Tasks: The performance of CTLE\$ drops by 59% in the Next POI Prediction (NPP) task and 60% in the Trajectory User Link (TUL) task, showing that the Transformer Encoder (TE) plays a crucial role in trajectory-related tasks.
Table 7. Results for JCLRNT variants on the Xi’an dataset.
Methods JCLRNT JCLRNTJCLRNT& JCLRNT* JCLRNT\$
Encoder Core Arc TE TE TE TE LSTM
GAT GAT GAT GAT GAT
Pretraining Task TokRI
AGCL
ATrCL
ASI MAE↓ 5.660 5.708 5.744 5.601 5.702
RMSE↓ 11.02 11.06 11.11 10.90 11.00
TTE MAE↓ 115.9 118.3 116.5 115.3 124.5
RMSE↓ 183.5 185.5 183.8 182.3 189.8
STS ACC@3↑ 0.797 0.673 0.702 0.809 0.782
MR↓ 4.746 12.30 5.828 3.176 5.325
Per Rank 2 5 4 1 3
Comparison of JCLRNT Variants. Among road segment MapRL models, JCLRNT achieves the best performance. As shown in Tab. 7, we experiment with the following JCLRNT variants. (a) JCLRNT: Maintains the original model architecture. (b) JCLRNT%: Removes the ATrCL pre-training task. (c) JCLRNT&: Removes the Augmented Graph Contrastive Learning (AGCL) pre-training task. (d) JCLRNT*: Adds a new Token Relation Inference (TokRI) pre-training task, using road type classification, as implemented in HRoad. (e) JCLRNT\$: Replaces the Transformer with an LSTM as the sequence encoder.
The performance results on downstream tasks are presented in Tab. 7. From the experimental results, we draw the following observations: $i$) Comprehensive Pre-training Boosts Performance: JCLRNT* achieves the best performance across all downstream tasks by incorporating three pre-training tasks—TokRI, AGCL, and ATrCL—each aligned with the token-based, graph-based, and sequence-based encoders, respectively. This suggests that each encoder type benefits from having a tailored pre-training task. $ii$) ATrCL Provides Richer Representations: JCLRNTunderperforms compared to JCLRNT& in trajectory-related tasks like Travel Time Estimation (TTE) and Similarity Trajectory Search (STS). This indicates that ATrCL, a sequence-based pre-training task, offers richer representation capabilities than AGCL for these tasks. Even with a sequence encoder, the absence of ATrCL limits the model’s ability to capture dependency effectively, impacting performance. $iii$) Minor Impact of Replacing the Transformer Encoder: JCLRNT\$ shows relatively minor performance degradation when the Transformer Encoder is replaced by an LSTM. This suggests that the introduction of the graph encoder in JCLRNT helps extract relational information from network data, reducing the model’s reliance on the Transformer.
Table 8. Results from HREP variants on the Xi’an dataset.
Methods HREP HREPHREP& HREP* HREP\$
Encoder Core Arc HeteGCN HeteGCN HeteGCN HeteGCN GCN
Pretraining Task NFI
GAu
NCL
LPC ACC@1↑ 0.611 0.566 0.413 0.663 0.505
F1↑ 0.494 0.436 0.273 0.495 0.340
FI MAE↓ 127.9 128.1 133.4 119.7 146.6
RMSE↓ 312.3 320.8 323.7 269.8 325.3
MI MAE↓ 3.949 3.993 4.371 3.914 4.294
RMSE↓ 17.05 18.58 17.69 15.62 16.55
Per Rank 2 3 5 1 4
Comparison of HREP Variants. Among land parcel MapRL models, HREP achieves the best performance. As shown in Tab. 8, we experiment with the following HREP variants. (a) HREP: Maintains the original model architecture. (b) HREP%: Removes the Graph Autoencoder (GAu) pre-training task. (c) HREP&: Removes the Neighborhood Contrastive Learning (NCL) pre-training task. (d) HREP: Adds a new Node Feature Inference (NFI) pre-training task, incorporating human mobility inference, as implemented in MGFN. (e) HREP\$: Replaces the heterogeneous GCN (HeteGCN) with a standard GCN as the graph encoder.
The performance results on downstream tasks are presented in Tab. 8. From the experiments, we draw the following observations. $i$) NFI Enhances Flow Inference Performance: HREP* achieves the best results across all downstream tasks, including a 10% improvement in the Flow Inference (FI) task, indicating that the NFI pre-training task significantly enhances the model’s ability to predict flow-related information. $ii$) GAu Provides Richer Representations: HREPoutperforms HREP&, suggesting that the GAu pre-training task is crucial for capturing relationships between entities and improving downstream task performance. $iii$) HeteGCN Captures Diverse Entity Relationships: HREP\$ experiences a significant performance drop, especially in the Land Parcel Classification (LPC) task, when the HeteGCN is replaced with a GCN. This indicates that HeteGCN better captures the diverse relationships between different types of entities, making it essential for land parcel representation learning.
Summary for Variant Experiments. From the above experiments, we derive the following conclusions.
1) More Types of Pre-training Tasks Lead to Better Performance: Across all experiments, the best-performing models consistently incorporate the most diverse types of encoders and pre-training tasks. This demonstrates that integrating a wider range of pre-training tasks enhances the model’s ability.
2) Suitable Pre-training Tasks Provide Greater Benefits: Pre-training tasks tailored to specific downstream tasks significantly enhance performance. For instance, the Augmented Trajectory Contrastive Learning (ATrCL) task improves performance in trajectory-relatd tasks, such as Travel Time Estimation (TTE) and Similarity Trajectory Search (STS), while the Node Feature Inference (NFI) task greatly benefits Flow Inference (FI) tasks. This highlights the importance of aligning pre-training tasks with the specific requirements of downstream tasks to maximize model effectiveness.
3) Encoder Architecture Matters for Task Performance: The choice of encoder architecture has a significant effect. For instance, replacing the Transformer Encoder with an LSTM in CTLE and JCLRNT led to performance drops, especially in trajectory-related tasks. Similarly, HeteGCN outperforms a standard GCN in HREP by capturing more diverse entity relationships.

### v2

@@ 4.2.2. Experimental Results for Road Segment MapRL
2) Multiple Pre-training Tasks Improve Performance: Models incorporating multiple pre-training tasks, such as Toast, HRoad, START, and JCLRNT, consistently outperform those using a single pre-training task, like HRNR and SARN. Among token-based + graph-based models, HRoad, which integrates both NFI and GAu tasks, surpasses HRNR (GAu only) and SARN (AGCL only), highlighting the benefits of leveraging diverse pre-training tasks. By extracting information from multiple perspectives, these models learn more comprehensive feature representations. Additionally, in full-pipeline models, JCLRNT, which combines a sequence-based task (Augmented Trajectory Contrastive Learning) with a graph-based task (Augmented Graph Contrastive Learning), outperforms START, which employs two sequence-based tasks (MTR and ATrCL). This suggests that even among multi-task models, greater task diversity further enhances performance.
3) Fluctuating Results for SARN, Toast, START, and JCLRNT: These models experience higher performance variability compared to other models, largely due to their dependence on contrastive pretraining techniques such as Token Relation Contrastive Learning (TRCL), Augmented Graph Contrastive Learning (AGCL), and Augmented Trajectory Contrastive Learning (ATrCL). These methods are particularly vulnerable to shifts in training data splits and the inherent randomness in negative sample selection, leading to inconsistent results across multiple runs.
@@ 4.2.3. Experimental Results for Land Parcel MapRL
Table 5 presents the performance of land parcel MapRL models on the downstream tasks of Land Parcel Classification (LPC), Flow Inference (FI), and Mobility Inference (MI). The following observations can be made:
@@
3) Fluctuating Results for MGFN: Although MGFN ranks third overall (with an average rank of 3.53), its performance varies considerably under different random seeds, indicating a degree of instability. This instability arises because MGFN relies on clustering techniques to build the base relation network for representation learning, making it more sensitive to randomness.
4) Small Datasets Mitigate Instability in Negative Sampling: ZEMob, GMEL, and ReMVC are among the most stable models across all datasets. Despite employing contrastive learning as pretraining tasks—which tends to be sensitive in other MapRL models—their stability may be attributed to the smaller number of parcel entities. For instance, San Francisco contains 194 parcel entities compared to 27,274 road segment entities. This smaller number of parcel entities likely results in more consistent negative sampling. Additionally, tasks for land parcels are simpler, focusing on static attribute inference rather than complex sequential predictions, which further contributes to stable learning.
@@ 4.2.4. Summary for Overall Experiments
3) Role of Auxiliary Data: Incorporating trajectory, and relation network data improves model performance across tasks. However, models that fully exploit trajectory dependencies achieve diminishing gains from relation networks, indicating some redundancy between these data sources.
4) Instability Introduced by Contrastive Learning: Contrastive learning relies on the generation of positive and negative sample pairs, which vary between different random seeds. Consequently, models using contrastive learning pretraining tasks are more sensitive to randomness. However, this sensitivity may decrease when working with smaller datasets.
@@
4.3. Scalability Analysis of MapRL Models.
Scalability is a critical factor in MapRL model design, determining their adaptability to varying data sizes and computational constraints. In this section, we analyze scalability from three key perspectives: effectiveness, efficiency, and the trade-offs between model size and inference time. Effectiveness assesses how well models maintain performance as the amount of fine-tuning data varies. Efficiency evaluates computational costs, including parameter count, training time, and inference time, across datasets with different entity scales. Finally, the trade-off analysis explores how model complexity impacts inference speed, offering insights into balancing accuracy and deployment feasibility.
4.3.1. Effectiveness
In this experiment, we adjust the fine-tuning data size from 10% to 50% of the original training set and compare the result of pre-built models in VecCity. POI-oriented models are tested on the Tokyo dataset, while road segment- and land parcel-oriented models are evaluated on the Xi’an dataset. Fig. 3 presents the evaluation result: Fig. 3 (a) to (c) show results for POI-oriented models. Fig. 3 (d) to (f) display results for road-segment-oriented models. Fig. 3 (g) to (i) provide results for land-parcel-oriented models. From these figures, we make the following observations:
1) Performance Scalability. Token-based models (dashed lines) show limited scalability, with performance gains plateauing around 30% of fine-tuning data. In contrast, graph- and sequence-based models continue improving up to 50%. Notably, methods incorporating contrastive learning techniques—such as CACSR, START, JCLRNT, and HREP—show particularly strong improvements, suggesting that contrastive learning further enhances performance as more data becomes available.
@@ 4.3.1. Effectiveness
2) Stability in Attribute Inference Tasks: MapRL models exhibit stable performance in attribute inference tasks (POIC, ASI, LPC) regardless of training data size. Since these tasks primarily rely on entity attributes learned during pre-training, only a small amount of labeled data is needed for fine-tuning.
Figure 4. Scalability of MapRL models in model size, training time, and inference time across datasets with varying entity counts. Each point represents a dataset, with lines connecting them in ascending entity count to highlight trends.
3) Fluctuations in Trajectory-Related and Mobility Inference Tasks: Models exhibit notable performance fluctuations in trajectory-related tasks (e.g., NPP, TUL, TTE) and mobility inference tasks (e.g., FI, MI). Trajectory-related tasks require sequence-level predictions, whereas most encoders and pre-training tasks generate single-entity embeddings, making these tasks especially sensitive to smaller fine-tuning datasets. By contrast, Similarity Trajectory Search (STS) relies on similarity comparisons rather than prediction, making it more stable. Additionally, Flow Inference (FI) and Mobility Inference (MI) capture human mobility flows that vary greatly over time. This temporal heterogeneity demands more labeled data to accurately model these patterns, leading to further performance variability when data are limited.
@@
Overall, token-based models show weaker performance scalability, with diminishing returns as dataset size increases. In contrast, graph-based and sequence-based models, especially those utilizing contrastive learning methods (e.g., CACSR, START, JCLRNT, HREP), show greater performance improvements as labeled data increases, demonstrating stronger scalability in data-intensive scenarios. Moreover, existing MapRL models exhibit varying sensitivity to the size of labeled data across different tasks. For tasks less influenced by temporal information, such as attribute inference (ASI, POIC, LPC) and Similarity Trajectory Search (STS), models maintain stable performance even with limited training data. However, for tasks with time-dependent labels, such as trajectory-related tasks and mobility inference tasks, larger amounts of labeled data are necessary to effectively capture temporal dynamics.
4.3.2. Efficiency
We evaluate model’s efficiency across three key aspects: (1) Parameter count, representing the space complexity of the models; (2) Epoch time, indicating the time cost of training per epoch; and (3) Inference time, reflecting the time required for downstream task evaluation. These measurements are collected from all five datasets used in the main experiments, each with different entity counts. Figure. 4 presents the evaluation results: Fig. 4 (a) to (c) show results for POI-oriented models, Fig. 4 (d) to (f) display results for road segment-oriented models, and Fig. 4 (g) to (i) provide results for land parcel-oriented models. The horizontal axis of all figures represents the number of entities on a $10^{3}$ scale, while the vertical axis represents the parameter count in millions, time cost per epoch in seconds, and total inference time in seconds, respectively. Based on these figures, we have several key findings:
1) Parameter Count: Token-based models (dashed lines) generally achieve better scalability in model size, while sequence-based models (e.g., CACSR, Toast, START) often have higher parameter counts, likely due to the increased space complexity of sequence encoders. In contrast, SARN, GMEL, MVURE, and HREP show relatively stable parameter counts as the number of entities increases, demonstrating better scalability in scenarios with large numbers of map entities.
Figure 5. Tradeoff between model parameters and time efficiency on average, across all available datasets.
2) Training Time: Token-based models (dashed lines) demonstrate better training time scalability in POI-oriented models but perform less efficiently in land parcel-oriented models. This difference is influenced by entity count variations, with POI datasets containing $1\times 10^{4}$ to $8\times 10^{4}$ entities, whereas land parcel datasets range from $2\times 10^{2}$ to $1\times 10^{4}$. The results suggest that token-based models may have a greater competitive advantage in large-scale datasets. Sequence-based models (e.g., CACSR, Toast, START) generally incur longer training times. Graph-based models employing GAu pretraining tasks (HRNR, GMEL, MGFN) show more-than-linear growth in training time, posing scalability challenges for large datasets, whereas HREP, despite using GAu pretraining, maintains relatively stable training time due to its lower space complexity.
3) Inference Time: All models show an increasing trend in inference time. However, the rise is more pronounced in POI- and land parcel-oriented models compared to road segment-oriented models, suggesting that inference time is more sensitive to entity count in these tasks. Sequence-based models (e.g., CTLE, Toast, START, JCLRNT) generally require longer inference times. This can be attributed to their reliance on sequential encoders, which introduce additional computational steps.
4.3.3. Trade-offs.
In Section 4, we observe that complex models tend to achieve higher accuracy (e.g., full pipeline models). However, increased model complexity can lead to efficiency challenges, particularly in inference time. The previous efficiency analysis does not directly examine the trade-off between model complexity and inference time. To address this, we evaluate the relationship between parameter count (millions) and single-sample inference time (ms). This experiment uses the same dataset as in Section 4.3.1 on effectiveness: POI-oriented models are evaluated on the Tokyo dataset, while others are tested on the Xi’an dataset.
Fig. 5 presents the results, showing that land-parcel-oriented models maintain relatively stable inference times despite increases in model size, likely due to the lower complexity of their downstream tasks. In contrast, POI and road segment-oriented models exhibit a clear upward trend in inference time as model size grows. Sequence-based models (e.g., CTLE, JCLRNT, Toast, START) have higher model size and inference times due to the integration of sequence encoders. Meanwhile, graph-based models like MVURE, GMEL, ReMVC, and HREP generally maintain lower inference times and smaller model sizes. However, HRNR incurs higher inference costs due to its multi-level GNN encoder.
((a)) Top 3 combinations for POI map entity.
((b)) Top 3 combinations for segment map entity.
((c)) Top 3 combinations for parcel map entity.
Figure 6. Top 3 pretraining task combinations.
5. Delving into Pretraining Tasks.
In our experimental analysis in Section.4, we conclude that utilizing multiple pretraining tasks can enhance the performance of MapRL models. Moreover, greater task heterogeneity—incorporating token-based, graph-based, and sequence-based pretraining tasks—can further improve model capability. To fully investigate the influence of task heterogeneity on performance, we conduct deep dive experiments to figure out what specific combinations of heterogeneous pretraining tasks can yield the best gains on MapRL.
Experimental Space. In this section, we aim to identify the best pretraining task combinations for MapRL. The candidate combination set is constructed by combining different categories of pretraining tasks. For POI entities, we consider token-based pretraining tasks (TokRI, TRCL) and sequence-based pretraining tasks (TrajP, MTR, ATrCL), forming combinations such as TokRI+TrajP, TRCL+MTR, etc. Similarly, for road segments and land parcels, candidate sets are formed by combining token-, graph-, and sequence-based pretraining tasks (if applicable), as shown in Table 6.
Table 6. Available pretraining tasks by entity types.
Task Type POI Road Segment Land Parcel
Token-based TokRI, TRCL TokRI, TRCL TokRI, TRCL, AToCL
Graph-based NFI, GAu,AGCL NFI, GAu,NCL
Seq-based TrajP, MTR, ATrCL TrajP, MTR, ATrCL
The candidate set is constructed based on three rules: (1) selecting pretraining tasks implemented and validated by existing models; (2) using available datasets for testing; and (3) repeating all experiments five times with different random seeds. To ensure experimental feasibility, we further narrow the experimental scope by selecting only one pretraining task from each category for candidate combinations. The candidate set for other entity types follows the same principle.
Model Architecture. To eliminate the influence of model architecture, we use the same encoder within each type of MapRL model and include only the necessary encoding components. Specifically, token-based encoders map entity IDs to embeddings. Additionally, the token encoder for POI incorporates spatial features by transforming coordinates into a vector using an MLP, while the token encoder for land parcel integrates semantic features in a similar way. Graph-based encoders use GCNs to process relation networks. For land parcels, the relation network defines edges using POI similarity, spatial distance, and human mobility transitions, following HREP (Zhou et al., 2023). Sequence-based encoders leverage Transformer layers to encode sequential auxiliary data. The model is configured with the same setting as the main experiment in Section 4. Encoders are arranged in the order: Token → Graph → Sequence.
Experimental Result. To ensure a focused discussion, we highlight the top three pretraining task combinations for each entity type, ranked by their average performance across tasks and datasets. The results are shown in Fig. 6, where each subfigure corresponds to a specific downstream task or dataset. The horizontal axis represents rankings (lower is better), while the vertical axis lists the top three combinations, with lighter colors indicating lower ranks. In Fig. 6(a) for POI entities, the first three subfigures show each model’s average ranking in a specific downstream task (POIC, NPP, or TUL) across multiple datasets. We refer to these as task-oriented rankings. In contrast, the last five subfigures show each model’s average ranking across all tasks within a specific dataset (NY, TYO, CHI, SIN, or SF), which we refer to as dataset-oriented rankings. The same ranking logic applies to Fig. 6(b) for road segments and Fig. 6(c) for parcels. Based on the results, we have several key findings:
1) Optimal Combinations for POI: TokRI + MTR consistently ranks highest, likely due to its ability to model POI categories and long-range temporal dependencies—two key attributes for POIs. However, in SF, where data sparsity is more pronounced, TokRI + TrajP emerges as a more suitable alternative, as TrajP focuses on local temporal patterns, making it more reliable in sparse data scenarios.
2) Optimal Combinations for Road Segment: AGCL outperforms others in all scenarios, as it effectively captures important spatial dependencies in geographic relation (GR) networks, which is critical for modeling road segments. From a dataset-oriented perspective, AGCL + MTR achieves the best results in CD and PRT, where visit frequencies are higher (CD: 1551, PRT: 2430), allowing MTR to learn sequential dependencies from dense trajectory data. In contrast, AGCL + TrajP works better in BJ and SF, where visit frequencies are lower (BJ: 856, SF: 771). This suggests that MTR benefits more from dense data, while TrajP is better suited for sparse data.
3) Optimal Combinations for Land Parcel: AToCL achieves the best performance in 5/8 cases, likely because its augmentation technique enhances the model’s ability to capture key semantic features while filtering out noise. This is especially important for land parcels, where POI category distributions define functional attributes. GAu and NFI are also well-suited for land parcels, as they better model similarity-driven connectivity in social relation (SR) networks.
By integrating these insights, we provide a general guide for selecting pretraining tasks across different entity types:
1) Token-based: Encoding semantic features is a key capability of token-based encoders. TokRI is well-suited for structured categorical attributes, while AToCL’s augmentation-based contrastive learning better handles high-dimensional and redundant semantic features, making it particularly effective for land parcels.
2) Graph-based: AGCL is well-suited for geographic relation networks, as it effectively captures key spatial dependencies. In contrast, GAu and NFI perform better in social relation networks due to their ability to model similarity-driven connectivity.
3) Sequence-based: MTR excels at capturing long-range sequential dependencies, making it the best choice for sequence-based tasks, but it requires dense data to be effective. In contrast, TrajP is more reliable in sparse data scenarios.